OpenStudy (anonymous):

When you divide two fractions, is the quotient always smaller than the dividend? A) No. The quotient is sometimes larger. B) No. The quotient is always larger than the divisor. C) There is not enough information to answer the question. D) Yes. Division means that you are dividing a number into smaller pieces.

OpenStudy (kidrah69):

what do you think?

OpenStudy (anonymous):

A

OpenStudy (kidrah69):

When you divide fractions, like 1/4 divided by 1/4, the answer is a larger number than 1/4. so i think you are correct

OpenStudy (kidrah69):

Quotients can be greater at times

OpenStudy (amistre64):

d/n = q let d=a/b let n = x/y let q > d \[\frac{a/b}{x/y}>\frac ab\] \[\frac{ay}{bx}>\frac ab\] if x=1. and y>0, then this would satisfy that q > d 2/3 --- = 4/3 1/2 so A is fine

OpenStudy (amistre64):

as long as x<y, this would make q>d